A solution is prepared by dissolving 26.7 grams of sodium hydroxide in 650 grams of water. What is the mole fraction of the sodium hydroxide?
1
Expert's answer
2019-03-18T02:20:10-0400

c(NaOH)="\dfrac{m(NaOH)}{M(NaOH)*V}=\dfrac{26.7g}{40g/mol*0.65L} = 1.027 mol/L"